OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 8, x = 2. Find y when x = 8. A.2 B.16 C.24 D.32

OpenStudy (anonymous):

Inverse variation \[y=\frac{k}{x}\] Plug in your x and y to solve for k.

OpenStudy (anonymous):

Remember to replace x with 2x because "twice the value of x"

OpenStudy (cwrw238):

y varies as 1/2x means y = k/2x where k is some constant

OpenStudy (anonymous):

i keep getting B.

OpenStudy (anonymous):

That's k, you have to plug in the new x to find the new y.

OpenStudy (anonymous):

Take that value of k that you got, and put it and the new x=8 into the equation to solve for y.

OpenStudy (anonymous):

So 2?

OpenStudy (anonymous):

Oh, wait, what did you get for k? k=32, so yes, y = 2 at x=8 because 2=32/16.

OpenStudy (anonymous):

Yes, so i guess ill go with A. 2

OpenStudy (anonymous):

\[Inverse \space variation: y=\frac{k}{2x}, \space k=32, \space x=8, \space y=\frac{32}{16}.\]